OpenStudy (anonymous):

A garrison of 1500 men is provisioned for 60 days. After 25 days, the garrison is reinforced by 500 men. How long will the remaining provisions last?

OpenStudy (mathstudent55):

Each man needs x provisions per day. 1500 men need 1500x provisions in one day 1500 men need 1500x * 60 = 90,000x provisions in 60 days In 25 days, they consume 25 * 1500x = 37500x provisions After 25 days, they have left 90,000x - 37,500x = 52,500x provisions left 2000 men need 2000x provisions in one day 52,500x provisions divided by 2000x provisions per day is 26.25 days